OpenStudy (anonymous):

Convert These To Decimals!! :) 13/25= 43/100= 29/50=

OpenStudy (anonymous):

13/25 = .52 43/100 = .43 29/50 = .58

OpenStudy (anonymous):

13/25 = 52/100 = .52 43/100 = .43 29/50 = 58/100 = .58